Operation Summary for April 18, 2011 - April 20, 2011 |
|
Monday April 18 The day shift began with Operations monitoring the Tevatron (TEV) store 8666 (186.68E30) and theRecycler stash (103.7E10), with the Antiproton Source (Pbar) stacking (29.4E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTFB T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations increased the NuMI beam intensity to 1.6E13 protons per pulse. The TEV quenched at 2 PM due to the failure of a quench protection system control card. Operations sent beam to the MuCool (MTA) experiment. During the evening shift, Operations disabled beam to T-978 at 6 PM. Operations put the TEV into HEP shot setup at 8 PM and established store 8670 at 9:05 PM with an average initial luminosity (AIL) of 328E30.Tuesday April 19 The midnight shift began with Operations monitoring store 8670 (189.3E30) and the stash (104.5E10), and with Pbar stacking (10.8E10). During the midnight shift, Operations established beam to T-978 at 4:15 AM. During the day shift, Operations terminated store 8670 at 12:55 PM and put the TEV into shot setup. Operations abandoned the shot setup after antiproton transfers #2, #3, and #4 failed to load into the TEV. Operations tracked this problem to an errant Main Injector ramp file being loaded into MECAR by an MI studier. During the evening shift, Operations recovered the TEV, put it into shot setup at 5:15 PM, and established store 8672 at 6:40 PM with an AIL of 281E30. However, at 7:20 PM, a lightning strike aborted the store. There was no quench. MTA personnel began conducting studies at 7:40 PM. Operations recovered the TEV and TEV personnel began conducting electron cloud studies at 10 PM.Wednesday April 20 The midnight shift began with the TEV and MTA personnel conducting studies, with Operations monitoring the stash (175.9E10), with Pbar stacking (9.7E10), and with MiniBooNE taking beam. During the midnight shift, TEV and MTA personnel completed their studies. The FESS Duty Electrician switched off Feeder 23 so that operators and FESS personnel could access TEV sector A2 to investigate a sump pump alarm. FESS personnel reported that the sump pump motor was problematic. FESS personnel installed a temporary pump and operators switched the TEV back on. D0 personnel reported that they needed to access their collision hall. Operators halted the TEV checkout and prepared D0 for an access. CDF personnel took advantage of the downtime to access their collision hall. Operations established beam to T-978 at 4:50 AM. The Plans for Wednesday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 04/20/2011 email Fermilab |
| Security, Privacy, Legal | |